OpenStudy (sanic):

Remo needs to buy a countertop for a bathroom. He calculated the area as 10 square feet. The actual area is 9.5 square feet. What is Remo's percentage error? 0.5% 0.95% 5% 5.26%

OpenStudy (sanic):

I'll give medal

OpenStudy (sanic):

I'll give medal

OpenStudy (lochana):

\[percentage \ error = \frac{mesured \ value - actual\ value}{actual\ value}100\]

OpenStudy (sanic):

wait so what would the answer be

OpenStudy (anonymous):

\[\frac{(10-9.5) 100}{9.5}=5.26 \% \]

OpenStudy (sanic):

What

OpenStudy (sanic):

That wasn't an answer

OpenStudy (lochana):

we are not allowed to give direct answers. yes you already had it:)